# The Analytic Stockwell Transform and its Zeros

### v2

@@ 4.2 Spatial Statistics
Figure 3: Theoretical pair correlation function of the zeros of the hyperbolic Gaussian Analytic function of parameter $\alpha$ (dashed black line) and averaged empirical pair correlation functions (blue solid line) of the zeroes of the Analytic Stockwell Transform of parameter $\beta=(\alpha-1)/2$ of white noise for $\alpha\in\{100,300,500\}$. Ensemble averaging and quantile estimation are performed over $R=100$ realizations of the discrete white noise in dimension $N=4000$, corresponding to a time period of $x_{\max}-x_{\min}=1$ s long and a sampling frequency $\nu_{s}=4000$ Hz.
@@ 5 Conclusion and Perspectives
After evidencing the existence of an Analytic Stockwell Transform, whose analysis window have been designed leveraging a modulated Cauchy wavelet, the zeros of the transform of white noise are characterized. To that aim a well-suited construction of continuous white noise has been presented, enabling to establish a strong connection between the Analytic Stockwell Transform of white noise and the hyperbolic Gaussian Analytic Function. This link permitted to transfer the remarkable properties of the zeros of the hyperbolic Gaussian Analytic Function, among which the invariance under isometries of the Poincaré disk, to the zeros of the Analytic Stockwell Transform of white noise. Finally, intensive Monte Carlo simulations are implemented to illustrate and support this result; notably the empirical pair correlation function of the zeros of the Analytic Stockwell Transform of white noise is shown to coincide very accurately with the expected theoretical pair correlation function of the zero set of the hyperbolic Gaussian Analytic Function. A documented Python toolbox reproducing all the experiments has been produced and made publicly available. [^7] https://github.com/courbot/ast
The characterization of the distribution of zeros under a white noise hypothesis is the cornerstone of recent original zero-based detection procedures [5, 30, 27]. Hence, the direct continuation of the present work would consists in leveraging the knowledge gained on the zeros Analytic Stockwell Transform to perform signal detection in high-noise contexts. Though, application of spatial statistics in signal processing do not restrict to detection and zero-based denoising and separation strategies have shown promising results [15, 5, 27]. The nice resolution properties and great versatility of the Analytic Stockwell Transform make it a good candidate for the development of zero-based component separation algorithms. Finally, as the accuracy of zero-based procedures relies on the precise localization of zeros, the off-the-grid algorithm developed in [9] will be generalized to the hyperbolic geometry framework to enhance the quality of spatial statistics estimators, and hence reaching state-of-the-art detection and reconstruction performance.

### v3

@@ 4.2 Spatial Statistics
Figure 3: Comparison of theoretical and empirical spatial statistics of the zeroes of the Analytic Stockwell Transform of parameter $\beta=(\alpha-1)/2$ of white noise for $\alpha\in\{50,100,300,500\}$. (Left) First intensity $\rho_{1}^{(\alpha)}$. (Right) pair correlation function $g^{(\alpha)}$. Theoretical spatial statistics, (48) and (49), are represented as dashed black lines and averaged empirical statistics, (50) and (51), as solid blue lines. Ensemble averaging and quantiles estimation are performed over $R=100$ realizations of the discrete white noise of $N=4000$ points, corresponding to a time period of $x_{\max}-x_{\min}=1$ s long and a sampling frequency $\nu_{s}=4000$ Hz. The $x$-axis is expressed in pseudo-hyperbolic distances (13).
@@ 5 Conclusion and Perspectives
After evidencing the existence of an Analytic Stockwell Transform, whose analysis window have been designed leveraging a modulated Cauchy wavelet, the zeros of the Analytic Stockwell Transform of white noise are characterized. To that aim a well-suited construction of continuous white noise has been presented, enabling to establish a strong connection between the Analytic Stockwell Transform of white noise and the hyperbolic Gaussian analytic function. This link permitted to transfer the remarkable properties of the zeros of the hyperbolic Gaussian analytic function, among which the invariance under isometries of the Poincaré disk, to the zeros of the Analytic Stockwell Transform of white noise. Finally, intensive Monte Carlo simulations are implemented to illustrate and support this result; notably the empirical pair correlation function of the zeros of the Analytic Stockwell Transform of white noise is shown to coincide accurately with the expected theoretical pair correlation function of the zero set of the hyperbolic Gaussian analytic function. A documented Python toolbox reproducing all the experiments has been produced and made publicly available. [^9] https://github.com/courbot/ast
The characterization of the distribution of zeros under a white-noise hypothesis is the cornerstone of recent original zero-based detection procedures [6, 31, 28]. Hence, the direct continuation of the present work would consists in leveraging the knowledge gained on the zeros Analytic Stockwell Transform to perform signal detection in high-noise contexts. Though, application of spatial statistics in signal processing do not restrict to detection and zero-based denoising and separation strategies have shown promising results [16, 6, 28]. The nice resolution properties and great versatility of the Analytic Stockwell Transform make it a good candidate for the development of zero-based component separation algorithms. Although in this work, the phase information of the Stockwell Transform, which differs from the phase information contained in the Wavelet Transform, has not been exploited, it could interestingly be used to complete the information of zeros with the aim of consolidating and complementing methods for signal detection and denoising since it seems that both information are intimately related  [5]. Finally, as the accuracy of zero-based procedures relies on the precise localization of zeros, novel zero detection techniques will be explored. To go beyond the Minimal Grid Neighbor algorithm, two paths will be investigated: first, an hyperbolic counterpart of the Adaptive Minimal Grid Neighbor algorithm [13] will be derived, and second, the off-the-grid algorithm developed in [10] will be generalized to the hyperbolic geometry framework with the overarching goal of enhancing the quality of spatial statistics estimators, and hence reaching state-of-the-art detection and reconstruction performance.